OpenStudy (anonymous):

slope = -9/2 ,point (-2,-9), find equation of the line ?

OpenStudy (anonymous):

in order to determine the equation of the line , we use y-y1 = m(x-x1) where(x1,y1) is point on the line and m is the slope

OpenStudy (anonymous):

i know how to use this , but i confuse due to fraction..

OpenStudy (anonymous):

y - (-)9 = {-9/2}*{x-(-)2} y+9 = -9/2*(x+2) multiply both sides by 2 hope you can do now

OpenStudy (anonymous):

thanks for your help. gave you a medal for the question....